# Assessing Logical Reasoning Capabilities of Encoder-Only Transformer Models

### v1

@@ Abstract
Logical reasoning is central to complex human activities, such as thinking, debating, and planning; it is also a central component of many AI systems as well. In this paper, we investigate the extent to which encoder-only transformer language models (LMs) can reason according to logical rules. We ask whether those LMs can deduce theorems in propositional calculus and first-order logic; if their relative success in these problems reflects general logical capabilities; and which layers contribute the most to the task. First, we show for several encoder-only LMs that they can be trained, to a reasonable degree, to determine logical validity on various datasets. Next, by cross-probing fine-tuned models on these datasets, we show that LMs have difficulty in transferring their putative logical reasoning ability, which suggests that they may have learned dataset-specific features, instead of a general capability. Finally, we conduct a layerwise probing experiment, which shows that the hypothesis classification task is mostly solved through higher layers.
@@ 1 Introduction
Transformer models are remarkably effective at a wide range of natural language processing (NLP) tasks, such as question answering, summarization, and text generation. By and large, these abilities are the result of specific training processes, where a language model is fine-tuned on a task-specific dataset. Curiously, encoder-only transformer models (LMs) also exhibit implicit linguistic and cognitive abilities for which they were not directly supervised. Such LMs have been shown to encode information on tense and number (Conneau et al., 2018), anaphora and determiner-noun agreement (Warstadt et al., 2020), semantic roles (Tenney et al., 2019b), syntactic dependencies (Manning et al., 2020), relational knowledge (Dai et al., 2021), and spatiotemporal representation (Gurnee and Tegmark, 2023).
Hence, it is worth investigating whether encoder-only transformer models demonstrate logical reasoning capabilities as well. Given that logical reasoning plays a fundamental role in intelligence, human or otherwise (Russell and Norvig, 2010), it is essential to explore if these models exhibit such capabilities. Concerning NLP, understanding the structure of LMs’ logical reasoning may enable us to discern the extent to which their inferences and arguments arise from reasoning rather than purely associative memory. This understanding is crucial for developing mechanisms that facilitate the generation of factual text from non-structured databases.
@@
Section 2 reviews the work on transformers’ logical reasoning abilities, as well as the function of probing in uncovering latent knowledge. Next, we gather and describe four datasets grounded on logical deduction (sec. 3). In a first batch of experiments, we conduct a systematic comparison of encoder-only transformer models on these datasets (sec. 4). Section 5 then investigates whether the performance in the previous task could be attributed to some general ability and whether the reasoning learned from one dataset could be transferred to a similar dataset. Finally, we perform a layer-wise probing to understand which layers are responsible for solving logical deduction problems (sec. 6). [^1] Data, code, and complete results are available at https://github.com/paulopirozelli/logicalreasoning.
@@ 2 Related Work
Transformer models are powerful enough to solve logical reasoning tasks expressed in natural language (Clark et al., 2020; Zhang et al., 2022; Tian et al., 2021; Han et al., 2022). Yet, it is not clear if these models have actually mastered logical reasoning. LMs seem to inevitably rely on statistical artifacts to deduct theorems, rather than in general, rule-based relationships (or something akin to that) (Zhang et al., 2022). They use shortcuts to solve hypothesis classification problems, leading to vulnerabilities in reasoning (e.g., LMs are fooled when hypotheses appear within rules), and making them susceptible to irrelevant (logically-consistent) perturbations (Gaskell et al., 2022).
Studies focused on functional words close to logical operators have identified similar shortcomings in LMs’ reasoning capabilities. Transformers struggle to deal with negation, predicting similar probabilities to a sentence and its negation (Kassner and Schütze, 2020; Hossain et al., 2020). Kalouli et al. (2022) extend these findings to conjunction, disjunction, and existential and universal quantification, showing that expressions associated with these operations are frequently dominated by semantically rich words. Transformers also fail in modeling semi-functional-semi-content words in general, such as quantifiers and modals (Sevastjanova et al., 2021). Tangentially to logic, Dankers et al. (2022) show that transformer models do not always properly handle compositionality: sometimes a translation is more local than desired (treating idioms as regular constructions), sometimes it is excessively global (paying attention to irrelevant parts of the sentence). Moreover, there is evidence that the way LMs compose sentences does not align well with human judgment (Liu and Neubig, 2022).
Large LMs have also been assessed for their logical reasoning capabilities. Despite their impressive achievements in numerous tasks, these models still struggle with multi-step and compositional problems (Rae et al., 2021; Dziri et al., 2023). Although good at individual deductions, large LMs struggle with proof planing: when many valid proof steps are available, they often take the wrong path (Saparov and He, 2023). Large LMs also appear to suffer from human-like biases in logical tasks: they perform significantly worse when the semantic content is too abstract or conflict with prior knowledge (Dasgupta et al., 2022; Tang et al., 2023).
@@
It is usually hard to tell which data features are employed by an LM in solving a problem, and whether the LM is organizing its operations in human-like ways. By definition, fine-tuning is not a viable option to get insights about this process, as it requires reparametrizing the network; as a consequence, it is not possible to distinguish what has been learned in the task-oriented training and what was there from the beginning.
Probing tasks aim at recovering latent knowledge from the parameters of an LM: a dataset that encodes a particular property (e.g., POS) is used to train a classifier (the probe), taking the representations produced by the LM as inputs for the probe (Belinkov, 2022). As the LM is not further trained on the task, the performance of the probe depends on the information on that property been previously encoded by the model. Probing may be used to examine various components of LMs, such as embeddings, attentions or feedforward computations.
Transformer models have been extensively studied through probing (Belinkov, 2022). Most attention has been given to BERT, which gave rise to a large literature on the properties encoded by this model (Rogers et al., 2020). RoBERTa has also been studied in some detail through probing; e.g., what abilities that model learns during training (Zhang et al., 2021; Liu et al., 2021) and its knowledge of semantic structural information (Wu et al., 2021).
Figure 1: Examples of logical reasoning arguments. The argument on the top is from FOLIO, a manually-written dataset; the one in the bottom is from RuleTaker, a dataset that uses a semi-synthetic approach.
@@ 3 Logical Reasoning Datasets
Table 1: Main features of the logical reasoning datasets. FOL stands for first-order logic, PC for propositional calculus, and CI for conjunctive implication. Labels are: FOLIO (False, True, Unknown), LogicNLI (Contradiction, Entailment, Neutral, Paradox), RuleTaker (False, True), SimpleLogic (False, True). The average number of premises and the average number of words per argument is calculated for the training set. Appendix B shows the full label distribution.
Dataset Size (train/val/test) Scope Type Label Ave. Premises Ave. Words
FOLIO 1003 / - / 204 FOL Manual 3 5.23 64.01
@@
In assessing logical reasoning, we restricted our analysis to datasets related to propositional calculus (PC) and first-order logic (FOL); FOL being an extension of PC that includes predicates and quantifiers. These logical systems offer a powerful formalism that balances simplicity and expressivity in representing and reasoning about statements and relationships. They are suitable to capture a wide range of knowledge and to formalize a large number of domains.
@@
In addition to being expressable in PC or FOL, we selected datasets that satisfied three other properties: i) observations had to be self-contained and to require no prior knowledge; ii) sentences needed to have correspondent translations in both logical formalism and natural language; and iii) hypotheses had to be declarative sentences. The first property lets one decouple logic from background knowledge, in order to assess pure reasoning capabilities; this prevented the LMs’ from using their implicit knowledge base. For that reason, we only considered pure logical datasets. Thus, we did not take into account other forms of reasoning, such as scientific reasoning (e.g., Clark et al. 2018; Lai et al. 2017; Yu et al. 2020); mathematical reasoning (e.g., Cobbe et al. 2021; Patel et al. 2021; Hendrycks et al. 2021); counterfactual reasoning (e.g., Qin et al. 2019; Yang et al. 2020a; O’Neill et al. 2021); planning (e.g., Valmeekam et al. 2023); inductive reasoning (e.g., Sinha et al. 2019; Tang et al. 2023); and abductive reasoning (e.g., Tang et al. 2023). The second property, the translation in a logical formalism, is what allows one to determine unambiguously the logical relationship between premises and hypotheses, while the natural language counterpart allows us to probe the LMs. Hence, we excluded datasets that lacked natural language translations, such as LTL (Hahn et al., 2020). The last property excluded QA datasets as well (e.g., Liu et al. 2020; Talmor et al. 2019), as they required the understanding of several types of questions (e.g., why, what) that are entangled with semantic and contextual knowledge (e.g., that “Alice” is the name of a person).
@@
In the end, four logical reasoning datasets were selected: FOLIO (Han et al., 2022), LogicNLI (Tian et al., 2021), RuleTaker (Clark et al., 2020), and SimpleLogic (Zhang et al., 2022); some examples of arguments can be seen in Figure 1. These datasets cover a wide range of variations in terms of construction (manual, semi-synthetic, synthetic), alignment with common sense, linguistic variability, and scope (PC, full or partial FOL). For instance, FOLIO is human-written; LogicNLI and RuleTaker both use a template which is then manually edited; and SimpleLogic is fully synthetic. FOLIO and LogicNLI encompass the full spectrum of FOL; RuleTaker is expressable in FOL, but only covers negation and conjunctive implications; [^2] Conjunctive implications are arguments of the form ($fact.[\land fact.]^{*}rule.[\land rule.]^{*}\Rightarrow hypothesis$). and SimpleLogic is restricted to conjunctive implications in PC. As regarding the number of labels, RuleTaker and SimpleLogic have True and False labels; FOLIO includes an Unknown label; and LogicNLI admits a fourth possibility, Paradox, where both a hypothesis and its negation can be deduced from the premises.
[…]
@@ 4 Testing LMs for Hypothesis Classification
The performance of LMs in logical reasoning, and in particular in the datasets described in the previous section, has been studied in scattered experiments, without a clear unified context that allows direct comparison. Due to this, we fine-tuned a wide range of pretrained encoder-only LMs in those datasets, applying them to a hypothesis classification task; i.e., determining the logical relationship between a set of premises and a conclusion. These results represent a small yet valuable contribution to the existing literature.
To ensure comprehensive coverage, eight families of encoder-only models were assessed: DistilBERT (Sanh et al., 2020), BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), Longformer (Beltagy et al., 2020), DeBERTa (He et al., 2020), AlBERT (Lan et al., 2019), XLM-RoBERTa (Conneau et al., 2020), and XLNet (Yang et al., 2020b). The full list of models is displayed in the left column of Table 2.
Table 2: Accurary comparison among several encoder-only Language Models (LMs) for the hypothesis classification task across various datasets: FOLIO (FL), LogicNLI (LN), RuleTaker (RT), and SimpleLogic (SL). The table depicts different model types in the left column. Results for RoBERTa-large, the LM picked for subsequent analyses, are in bold. Best and worst scores for each dataset are highlighted in blue and red, respectively.
FL LN RT SL
@@
BERT-B 52.45 48.75 98.91 92.31
BERT-L 61.76 42.50 99.94 91.88
RoB-B 58.82 62.90 98.04 92.87
RoB-L 64.71 72.70 99.78 90.83
Longf-B 62.75 58.05 99.92 94.21
Longf-L 62.25 74.15 99.94 92.37
DB-XS 54.41 65.30 99.81 91.67
DB-S 53.43 59.65 95.23 93.23
DB-B 60.78 66.70 99.81 93.72
DB-L 64.71 84.70 99.97 93.72
DB-XL 62.25 25.00 99.99 49.96
DB-XXL 71.57 25.00 50.02 50.04
AlB-B 57.35 66.80 99.91 92.17
AlB-L 56.37 66.20 99.88 91.53
AlB-XL 38.73 65.10 99.97 90.05
AlB-XXL 56.86 93.90 99.94 92.24
XLM-B 55.88 45.20 97.64 91.74
XLM-L 62.75 65.45 99.95 91.74
XLNet-B 58.33 55.00 99.19 94.28
XLNet-L 58.33 71.40 99.86 92.94
Larg. class 35.29 25.00 50.02 50.03
To fine-tune our models, we used a linear layer with dimension (embedding-length, labels) and dropout of 0.5. As input for the classification head, we used the last hidden embedding of the [CLS] token, as usual for classification tasks in NLP. Sequences were padded and truncated to the maximum length allowed by each LM. Models were trained for up to 50 epochs, using early stopping with a patience of 5 epochs. We used Adam as our optimizer ($\beta$ = [0.9, 0.999] and no weight decay) and experimented with two different learning rates (1e-5 and 1e-6). Models were chosen based on the loss for the validation set, and results reported for the test set. The only exception is FOLIO, for which only the validation set is publicly available; results are thus reported for the latter. We report accuracy as the standard metric, as classes in the datasets are balanced. Table 2 displays the best result achieved in each case. The best and worst scores for each dataset are highlighted in blue and red, respectively. We also provide a largest class baseline for comparison. [^3] Complete results, including the losses for each set and the F1-score (micro, macro, and weighted) for the test set, can be found at the project’s Github page.
Results show that the LMs were able to classify hypotheses with reasonable success. All of them came close to solve RuleTaker, with all but DeBERTa-XL achieving an accuracy in SimpleLogic above 90%. Results were comparatively lower for FOLIO and LogicNLI, but still LMs considerably surpassed the largest class baselines. A weaker performance observed in these datasets was expected, given their greater language variability and broader logical scope; and in the case of FOLIO, smaller size as well. Overall, the encoder-only transformer models worked well as soft reasoners (Clark et al., 2020), being able to successfully deduce theorems from premises expressed in natural language. Noteworthy exceptions were the performances of AlBERT-XL in FOLIO; DeBERTa-XL and -XXL in LogicNLI and SimpleLogic; and DeBERTa-XXL in RuleTaker.

### v2

@@ Abstract
Transformer models have shown impressive abilities in natural language tasks such as text generation and question answering. Still, it is not clear whether these models can successfully conduct a rule-guided task such as logical reasoning. In this paper, we investigate the extent to which encoder-only transformer language models (LMs) can reason according to logical rules. We ask whether these LMs can deduce theorems in propositional calculus and first-order logic, if their relative success in these problems reflects general logical capabilities, and which layers contribute the most to the task. First, we show for several encoder-only LMs that they can be trained, to a reasonable degree, to determine logical validity on various datasets. Next, by cross-probing fine-tuned models on these datasets, we show that LMs have difficulty in transferring their putative logical reasoning ability, which suggests that they may have learned dataset-specific features instead of a general capability. Finally, we conduct a layerwise probing experiment, which shows that the hypothesis classification task is mostly solved through higher layers.
Keywords:
Logical reasoning Language models Transformer Probing
@@ 1 Introduction
Transformer models are remarkably effective at a wide range of natural language processing (NLP) tasks, such as question answering, summarization, and text generation. By and large, these abilities are the result of specific training processes, where a language model is fine-tuned on a task-specific dataset. Curiously, encoder-only transformer models (LMs) also exhibit implicit linguistic and cognitive abilities for which they were not directly supervised. Such LMs have been shown to encode information on tense and number [7], anaphora and determiner-noun agreement [49], semantic roles [46], syntactic dependencies [32], relational knowledge [8], and spatiotemporal representation [15].
Given that logical reasoning is a core component of intelligence, human or otherwise [38], it is worth investigating the capabilities of encoder-only transformer models in executing tasks that necessitate such reasoning. Understanding whether LMs can solve logical problems, and the manner in which they tackle these problems, may enable us to understand the extent to which their inferences arise from reasoning rather than purely associative memory. This understanding is crucial for developing mechanisms that facilitate the generation of consistent outputs, whether in a neural-symbolic fashion or through the improvement of model architectures.
@@
Section 2 reviews the work on transformers’ logical reasoning abilities, as well as the function of probing in uncovering latent knowledge. Next, we gather and describe four datasets grounded on logical deduction (sec. 3). In a first batch of experiments, we conduct a systematic comparison of encoder-only transformer models on these datasets (sec. 4). Section 5 then investigates whether the performance in the previous task could be attributed to some general ability and whether the reasoning learned from one dataset could be transferred to a similar dataset. Finally, we perform a layerwise probing to understand which layers are responsible for solving logical deduction problems (sec. 6). [^1] Data, code, and complete results are available at github.com/paulopirozelli/logicalreasoning.
@@ 2 Related Work
Transformer models are powerful enough to solve logical reasoning tasks expressed in natural language [4, 55, 47, 17]. Yet, it is not clear if these models have actually mastered logical reasoning. LMs seem to inevitably rely on statistical artifacts to deduce theorems, rather than on general, rule-based relationships [55]. They use shortcuts to solve hypothesis classification problems, leading to vulnerabilities in reasoning (e.g., LMs are fooled when hypotheses appear within rules), and making them susceptible to irrelevant (logically consistent) perturbations [14].
Studies focused on functional words close to logical operators have identified similar shortcomings in LMs’ reasoning capabilities. Transformers struggle to deal with negation, predicting similar probabilities to a sentence and its negation [25, 22]. [24] extend these findings to conjunction, disjunction, and existential and universal quantification, showing that expressions associated with these operations are frequently dominated by semantically rich words. Transformers also fail in modeling semi-functional-semi-content words in general, such as quantifiers and modals [41]. Tangentially to logic, [9] show that transformer models do not always properly handle compositionality: sometimes a translation is more local than desired (treating idioms as regular constructions), sometimes it is excessively global (paying attention to irrelevant parts of the sentence). Moreover, there is evidence that the way LMs compose sentences does not align well with human judgment [28].
Large LMs have also been assessed for their logical reasoning capabilities. Despite their impressive achievements in numerous tasks, these models still struggle with multi-step and compositional problems [36, 12]. Although good at individual deductions, large LMs struggle with proof planing: when many valid proof steps are available, they often take the wrong path [40]. Large LMs also appear to suffer from human-like biases in logical tasks: they perform significantly worse when the semantic content is too abstract or conflicts with prior knowledge [10, 44].
@@
Probing is a technique used to discover if a model has acquired certain type of knowledge. In probing, a dataset that encodes a particular property (e.g., part-of-speech) is used to train a classifier (the probe), taking the representations produced by the original model as inputs for the classification [1]. As the LM is not further trained on the task, as it would be in fine-tuning, the probe’s performance depends on whether the information about that property had already been encoded by the model. Thus, success in the task provides some evidence that the model has stored such knowledge in its parameters. Probing can be used to examine several components of LMs, such as embeddings, attention heads, and feedforward computations.
Transformer models have been extensively studied through probing [1]. Most attention has been given to BERT, which gave rise to a large literature on the properties encoded by this model [37]. RoBERTa has also been studied in some detail through probing; e.g., what abilities that model learns during training [56, 30] and its knowledge of semantic structural information [51].
Figure 1: Examples of logical reasoning arguments. The argument at the top is from FOLIO, a manually-written dataset; the one at the bottom is from RuleTaker, a dataset that uses a semi-synthetic approach.
@@ 3 Logical Reasoning Datasets
Table 1: Main features of the logical reasoning datasets. FOL stands for first-order logic, PC for propositional calculus, and CI for conjunctive implication. The labels in the datasets are as follows: FOLIO (False, True, Unknown), LogicNLI (Contradiction, Entailment, Neutral, Paradox), RuleTaker (False, True), SimpleLogic (False, True). The average number of premises and the average number of words per argument refer to the training set statistics. Appendix 0.B shows the full dataset label distribution.
Dataset Size (train/val/test) Scope Type Label Avg. Premises Avg. Words
FOLIO 1003 / 204 / - FOL Manual 3 5.23 64.01
@@
In assessing logical reasoning, we restricted our analysis to datasets related to propositional calculus (PC) and first-order logic (FOL), with FOL being an extension of PC that includes predicates and quantifiers. These logical systems offer a powerful formalism that balances simplicity and expressivity in representing and reasoning about statements and relationships. They are suitable for capturing a wide range of knowledge and formalizing a large number of domains.
@@
In addition to being expressible in PC or FOL, we selected datasets that satisfied three other properties: i) observations had to be as self-contained as possible; ii) sentences needed to have corresponding translations in both logical formalism and natural language; and iii) hypotheses had to be declarative sentences. The first property aims to decouple logic from background knowledge in order to assess pure reasoning capabilities. While using natural language examples means implicit knowledge can never be completely erased, we opted for datasets that minimized this by explicitly stating prior knowledge or by using inference patterns where resorting to prior knowledge is unnecessary. For this reason, we only considered pure logical datasets and did not include other forms of reasoning such as scientific reasoning [3, 26, 54], mathematical reasoning [5, 34, 20], counterfactual reasoning [35, 52, 33], planning [48], inductive reasoning [42, 44], and abductive reasoning [44]. The second property, the translation into logical formalism, allows one to determine unambiguously the logical relationship between premises and hypotheses, while the natural language counterpart allows us to probe the LMs. Hence, we excluded datasets that lacked natural language translations, such as LTL [16]. The last property excluded QA datasets [29, 43], as they required the understanding of several types of questions (e.g., who, where) that are entangled with semantic and contextual knowledge (e.g., that “Alice” is the name of a person).
@@
In the end, four logical reasoning datasets were selected: FOLIO [17], LogicNLI [47], RuleTaker [4], and SimpleLogic [55]. Examples of arguments can be seen in Figure 1. These datasets cover a wide range of variations in terms of construction (manual, semi-synthetic, synthetic), alignment with common sense, linguistic variability, and scope (PC, full or partial FOL). For instance, FOLIO is human-written; LogicNLI and RuleTaker both use a template which is then manually edited; and SimpleLogic is fully synthetic. FOLIO and LogicNLI encompass the full spectrum of FOL; RuleTaker is expressible in FOL but only covers negation and conjunctive implications, [^2] Conjunctive implications are arguments of the form ($fact.[\land fact.]^{*}rule.[\land rule.]^{*}\Rightarrow hypothesis$). and SimpleLogic is restricted to conjunctive implications in PC. Regarding the number of labels, RuleTaker and SimpleLogic have True and False labels; FOLIO includes an Unknown label; and LogicNLI admits a fourth possibility, Paradox, where both a hypothesis and its negation can be deduced from the premises. Table 1 summarizes the main statistics of the datasets. Appendix 0.A provides a detailed description of the four datasets.
For all datasets, inputs are formatted as “fact1. fact2. ... factn. rule1. rule2. ... rulem <SEP> hypothesis.”, for which a label must be predicted. The output is simply a probability distribution over the possible labels, which varies from dataset to dataset.
@@ 4 Testing LMs for Hypothesis Classification
The performance of LMs in logical reasoning, including in the datasets described in the previous section, has been studied in scattered experiments without a clear unified context that allows direct comparison. Due to this, we decided to fine-tune a wide range of pretrained encoder-only transformer models on those datasets. [^3] We opted to explore encoder-only models because this type of architecture is well-suited for classification tasks. These models have access to the whole input sequence and are typically trained on discriminative tasks, such as masked language modeling. We modeled this problem as a hypothesis classification task, where the goal is to determine the logical relationship between a set of premises and a conclusion. To ensure comprehensive coverage, eight families of LMs were assessed: DistilBERT [39], BERT [11], RoBERTa [31], Longformer [2], DeBERTa [19], AlBERT [27], XLM-RoBERTa [6], and XLNet [53]. The full list of models is displayed in the left column of Table 2.
To fine-tune our models, we used a classification head with a single linear layer of dimension (embedding-length, labels), and applied a dropout of 0.5 to the linear operation. As input for the classification head, we used the last hidden embedding of the [CLS] token, as is customary for classification tasks in NLP. Sequences were padded and truncated to the maximum length allowed by each LM. Models were trained for up to 50 epochs, using early stopping with a patience of 5 epochs. We used Adam as our optimizer (with $\beta$ = [0.9, 0.999] and no weight decay) and experimented with two different learning rates (1e-5 and 1e-6). Models were selected based on the loss for the validation set, with results reported for the test set. The only exception is FOLIO, where only the validation set is publicly available; results are thus reported for this set. We report accuracy as the standard metric, as classes in the datasets are balanced. Table 2 displays the best result achieved in each case. The best and worst scores for each dataset are highlighted in blue and red, respectively. We also provide a largest class baseline for comparison.
Table 2: Accuracy comparison among several encoder-only transformer models for the hypothesis classification task across four datasets: FOLIO, LogicNLI, RuleTaker, and SimpleLogic. The models are listed in the left column of the table. The best and worst scores for each dataset are highlighted in blue and red, respectively. “Largest class” refers to the accuracy achieved by always selecting the class with the highest frequency in the training set. Results for RoBERTa-large, selected for subsequent analyses, are in bold.
FOLIO LogicNLI RuleTaker SimpleLogic
@@
BERT base 52.45 48.75 98.91 92.31
BERT large 61.76 42.50 99.94 91.88
RoBERTa base 58.82 62.90 98.04 92.87
RoBERTa large 64.71 72.70 99.78 90.83
Longformer base 62.75 58.05 99.92 94.21
Longformer large 62.25 74.15 99.94 92.37
DeBERTa xsmall 54.41 65.30 99.81 91.67
DeBERTa small 53.43 59.65 95.23 93.23
DeBERTa base 60.78 66.70 99.81 93.72
DeBERTa large 64.71 84.70 99.97 93.72
DeBERTa xlarge 62.25 25.00 99.99 49.96
DeBERTa xxlarge 71.57 25.00 50.02 50.04
ALBERT base 57.35 66.80 99.91 92.17
ALBERT large 56.37 66.20 99.88 91.53
ALBERT xlarge 38.73 65.10 99.97 90.05
ALBERT xxlarge 56.86 93.90 99.94 92.24
XLM-RoBERTa base 55.88 45.20 97.64 91.74
XLM-RoBERTa large 62.75 65.45 99.95 91.74
XLNet base 58.33 55.00 99.19 94.28
XLNet large 58.33 71.40 99.86 92.94
Largest class 35.29 25.00 50.02 50.03
Results show that the LMs were able to classify hypotheses with reasonable success. Almost all models came close to solving RuleTaker and achieved an accuracy above 90% in SimpleLogic. Results were comparatively lower for FOLIO and LogicNLI, but the LMs generally surpassed the largest class baselines by a considerable margin. A weaker performance observed in these two datasets was expected, given their greater language variability and broader logical scope; and in the case of FOLIO, its smaller size as well. Overall, the encoder-only transformer models worked relatively well as soft reasoners [4], being able to successfully deduce theorems from premises expressed in natural language. Noteworthy exceptions were the performances of AlBERT-XL in FOLIO, DeBERTa-XL and -XXL in LogicNLI and SimpleLogic, DeBERTa-XXL in RuleTaker, and the overall lower performance of DistilBERT.